OpenStudy (anonymous):

Simplify the expression. (csc^2xsec^2x)/(sec^2x+csc^2x)

OpenStudy (anonymous):

do we put everyting in terms of sin and cos?

OpenStudy (yrelhan4):

Yup. do that and you would be able to simplify it.

OpenStudy (anonymous):

I get to that point and IDK what to do though

OpenStudy (anonymous):

I usually do. You don't have to but I find that easiest.

OpenStudy (anonymous):

What do you have?

OpenStudy (anonymous):

(1/sin^2x)(1/cos^2x) all divided by 1/cos^2x + 1/sin^2x

OpenStudy (yrelhan4):

Now leave the numertor as it is. Simplify the denominator first.

OpenStudy (anonymous):

Ok now combine the two fractions in the denominator

OpenStudy (anonymous):

1/cos^2x+sin^2x=1/1? os 1 on the denominator?

OpenStudy (yrelhan4):

You'll take the LCM right? so you'll get sin^2 x + cos^2 x/(sin^2*cos^2 x)..

OpenStudy (anonymous):

no im confused...

OpenStudy (anonymous):

well the denominator was ...\[\frac{cos^2x+sin^2x}{sin^2xcos^2}\]

OpenStudy (anonymous):

how?

OpenStudy (yrelhan4):

how would you solve this.. 1/2 + 1/3?

OpenStudy (anonymous):

LCM but im confused i though the denominator was 1/cos^2x+sin^2x

OpenStudy (yrelhan4):

you're getting the second term wrong..cosec^2 x= 1/sin^2 x

OpenStudy (anonymous):

thats what I have

OpenStudy (anonymous):

ok i understand...you have to multiply each fraction by the other fractions cos or sin

OpenStudy (anonymous):

so they get the samedenominator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then for the denominator you have 1/sin^2cos^2

OpenStudy (anonymous):

is the answer 1?

OpenStudy (yrelhan4):

Exactly. so you get sec^4 x* cosec^4x Yeah? I skipped a step.

OpenStudy (anonymous):

you have to rationalize the denominator, so you multiply by sin^2xcos^2x on top and bottom and then everything multiplies and cancels out on top?

OpenStudy (yrelhan4):

Oh dang. yeah i forgot we had sin and cos in the denominator. Well done.

OpenStudy (anonymous):

Yes that is correct the final answer is 1! Nice job!

OpenStudy (anonymous):

So im right??? :D YAY! Thanks!